The city subway advances at 72k m/h and sounds a 330 Hz horn when passing through an intersection where there is a stationary observer. What are the frequencies of the sound the observer hears when the metro passes by?


1
Expert's answer
2021-10-24T18:28:17-0400

72 km/h is 20 m/s. The frequency heard by the observer is


"f=f_0\frac{c}{c-v_s}=330\frac{343}{343-20}=350\text{ Hz}."

"v_s" here is the speed of the source of sound (train).